Two steamers leave a port on a river at the same time, each is traveling 22 mph. HOw long will it take for them to be 110 mi apart

Answer by mananth(16949)

they are travelling away from each other
so add up their speeds
72+72=144 mph
d=110 miles
t=d/r
t=110/144=0.76 hours
0.76*60 =45.8 minutes